OpenStudy (anonymous):

Solve, using the substitution method. 4x + 2y = 11 x – 2 = – 2y A.3,1/2 B.0,-2 C.o,-11/2 D.7,-5/2

OpenStudy (anonymous):

i believe you subtract 4x and 4x from 2y

OpenStudy (anonymous):

i need help

OpenStudy (studygurl14):

4x + 2y = 11 x -2 = -2y --> x = 2 - 2y Substitute into the first equation 4(2 - 2y) + 2y = 11 8 - 8y + 2y = 11 Can you do the rest? @bryson1

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

u sure

OpenStudy (anonymous):

MATH GENIUS DID YOU GET THAT ONE OR NO? I KNOW I GOT IT TOO 

OpenStudy (anonymous):

i think i got it......... wait, nope sorry

OpenStudy (anonymous):

\[4x+2y=11 4(2-2y)+2y=11. \]

OpenStudy (anonymous):

not 114 sorry